RC Update and Run Plan for Aug 7, 2020
- 1 RC Meeting Information
- 2 General updates
- 3 Current Run Plan
- 4 Notices in Effect for Shift Crew
- 5 Planned Accelerator Configuration Changes
- 6 Useful Links
RC Meeting Information
RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room
Friday Aug 7 2020.
- Hall in beam permit. Recovering target polarization.
- Transformer blown last night bringing down 1L07: 
- Jay is anticipating optics issues when the machine comes back up.
- Still working to bring up the machine (1L07 seems to be coming up about now as of 4pm).
- Target calibration polarization work 
- Alarming coil current issue: , Investigated by Bill: 
- Went to controlled access to check on HMS Q1 vacuum , 
- Need higher current beam.
Recent polarization measurements
- 25% at 2pm 
Current Run Plan
Normal Production running
- Make sure operational restrictions are followed.
- Make sure the raster size is correct (4-4.5)
- Monitor beam parameters/configuration very often and post an entry like this at the beginning of your shift and when run conditions have changed
- Running on cell targets: raster always ON (4-4.5, size will be determined by the scraping checks), orbit locks ON, current ramp rate 1 muA/sec
- Always take 120Hz of EDTM
- Fill the BTA with realistic numbers. If the desired beam current is not delivered prorate your ABUs!
- Change the IHWP every 4 hours when running on glass cells
- Run plan
- Re-establish beam .
- If MCC can deliver up to 42 muA we will do a BCM calibration 
- Notify RC so we can start Moller run (poc: Bill Henry and Dave G.)
- Production on 3He; if not able to reach 30 uA, pro-rate the ABUs
- We will start with a 4mm raster and 3uA. Check for scraping. Stop run. Increase to 4.5mm raster and 3uA. Check for scraping. If ok, then increase to 5uA and check for scraping. Goal is to establish largest raster size with no scraping. We are making this check due to 1mm shift of 3He cell.
Notes for tomorrow's 7:45 meeting
- Planning for Simona's SoLID Cerenkov access on Tuesday, will need to coordinate to rotate HMS to 29 deg, crane work, etc. ~4-5 hours work
- Target polarization calibration and studies on Tuesday-Wednesday during beam off and Moller measurement
- Tgt people: swap EPR photodiode on Tuesday?
- Moller measurement on Wednesday
- After ~ 10 days of running Hall C will go to 1 pass for ~3 days
- None planned right now
When there is a down:
- If the down is 2+ hours call Simona: she will check with the RC to see if there is an opportunity for work in the Hall (SoLID Cherenkov parasitic test)
Notices in Effect for Shift Crew
- Beam positions: 
- Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
- Magnet policies
- Spectrometer GUI (
go_magnets) should always be visible on Shift Leader computer, hcdesk3
- HMS Q1 must be ramped down in 50A steps in all cases
- Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
- If any of the Spectrometers magnets trip, take a screenshot of the PSU window and check the hall for smoke or water with the camera before resetting the interlock and post it to hclog.
- Spectrometer GUI (
- Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
- Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
- Ramp rate lock instructions are here and added to checklist
Planned Accelerator Configuration Changes
- Bluejeans remote connection information:
To join the Meeting: https://bluejeans.com/861439920 To join via Room System: Video Conferencing System: bjn.vc -or-220.127.116.11 Meeting ID : 861439920 To join via phone : 1) Dial: +1.408.740.7256 (US (San Jose)) +1.888.240.2560 (US Toll Free)) +1.408.317.9253 (US (Primary, San Jose)) (see all numbers - http://bluejeans.com/numbers) 2) Enter Conference ID : 861439920